The density ρ and pressure P of a gas are related by the expression c=√P/ρ


Find the basic unit of c.


1
Expert's answer
2021-08-02T08:41:07-0400

The units for density are "\dfrac{kg}{m^3}", the units for pressure are "\dfrac{kg}{s^2\cdot m}". Thus, the units of "c" are:


"\sqrt{\dfrac{\dfrac{kg}{s^2\cdot m}}{\dfrac{kg}{m^3}}} = \sqrt{\dfrac{kg\cdot m^3}{s^2\cdot m\cdot kg}} = \dfrac{m}{s}"

Answer. m/s.